OpenStudy (anonymous):

Do I just multiply? The length of the side of a rectangular desk is given by 7x-9ycm. The length of front is given by 7x+9y cm. What is the algebraic expression for the area of the top of the rectangular desk? Include units as part of your answer.

OpenStudy (anonymous):

Area= (7x-9)(7x-9)=49x^2-81 cm^2

OpenStudy (anonymous):

a=l*w a=(7x-9y)(7x+9y) use FOIL method: tyty doncatch thought so, just needed confirmation

OpenStudy (anonymous):

yes, yes .. u got that..